# Magnetar field dynamics shaped by chiral anomalies and helicity

### v1

@@ Abstract
The chiral magnetic effect (CME)—a macroscopic manifestation of the quantum chiral anomaly—induces currents along magnetic field lines, facilitating mutual conversion between chiral asymmetry and magnetic helicity. Although the finite electron mass suppresses chiral asymmetry through spin-flip processes, we demonstrate that the CME remains effective and plays a significant role in the magnetic field evolution of magnetars. The magnetic helicity acts as a persistent internal source of chiral asymmetry, which mediates the redistribution of magnetic energy across spatial scales, without requiring an external energy source. Focusing on the neutron star crust, we show that this mechanism reshapes the magnetic field configuration inherited at birth and, within a hundred years, amplifies both toroidal and poloidal large-scale dipolar components (relevant for spin-down) up to $10^{14}$ G at the expense of small-scale structures. Our results offer a microphysical mechanism—alternative to traditional hydrodynamic dynamo models—establishing a new framework for understanding magnetar field dynamics.
@@
Beyond macroscopic hydrodynamic processes such as dynamos and turbulence, quantum field theory establishes a fundamental microscopic connection between chirality and helicity through the chiral anomaly. This phenomenon facilitates a two-way transfer between fermionic chiral asymmetry and magnetic helicity. Recent research has increasingly focused on linking the generation of large-scale magnetic fields to chiral asymmetries produced during core-collapse supernovae and the proto-NS phase [8, 9]. However, significant challenges persist, most notably the efficiency of spin-flip scattering processes induced by the finite electron mass, raising doubts about the CME’s relevance in these environments. The rapid, temperature-dependent spin-flip reactions reduce the chiral imbalance, constraining the effectiveness of the chiral instability mechanism [10, 11].
@@
NSs exhibit a complex internal structure with multiple fluid components, each possessing distinct hydrodynamical velocities. In the outer crust, a rigid ionic lattice severely limits ion mobility, allowing only electrons to flow freely and sustain the currents essential for magnetic field evolution. The inner crust introduces added complexity with superfluid neutrons, which partially decouple from the nuclear lattice and act as a neutral fluid component. Complexity peaks in the NS core, where coexisting superfluid neutrons and superconducting protons necessitate a sophisticated multi-fluid framework. Given this complexity, magnetic field evolution is typically studied using region-specific approximations. Here, we focus on the crust, where higher electrical resistivity compared to the core leads to shorter magnetic field evolution timescales—making crustal physics a key region for understanding magnetar activity (ages $\lesssim 10\,\mathrm{kyr}$). In this regime, where only (slowly drifting) electrons are mobile due to the solid ionic lattice, the Hall-MHD (or e-MHD) approximation applies [19]. Magnetic field evolution is then governed by the induction equation derived from Faraday’s law:
@@
Magnetic field evolution must be coupled to the evolution equation for the chiral number density $n_{5}\equiv n_{R}-n_{L}$ [10, 20], which includes both source and sink terms:
@@
$\frac{\partial(\bm{A}\cdot\bm{B})}{\partial t}$ $=$ $-2c\bm{E}\cdot\bm{B}-c\bm{\nabla}\cdot\left(\bm{E}\times\bm{A}\right),$ (5)
@@
[…]
@@
Since $\sigma_{e}$ depends on temperature, so does $\Gamma_{f}$, underscoring the need to consider the coupled evolution of temperature and magnetic field in simulations.
@@
$n_{5}(t)=\left(\frac{2\alpha}{\pi\hbar}\frac{\bm{E}\cdot\bm{B}}{\Gamma_{f}}~{}.\right)\left(1-e^{-\Gamma_{f}t}\right)+n_{5}^{0}e^{-\Gamma_{f}t}.$ (8)
@@
$n_{5}(t)\approx\frac{2\alpha}{\pi\hbar}\frac{\bm{E}\cdot\bm{B}}{\Gamma_{f}}$ (9)
@@
Equation (10) highlights the two main contributions to the chiral asymmetry: the numerator originates from magnetic helicity and increases when ×$\bm{\nabla}\times\bm{B}$ aligns with $\bm{B}$. This alignment characterizes the so-called helical magnetic fields, where toroidal and poloidal components reinforce each other. In contrast, the spin-flip term in the denominator suppresses chiral asymmetry. If the magnetic field configuration possesses a non-zero average magnetic helicity, a small but appreciable chiral asymmetry can be sustained despite the action of spin-flip processes.
We define the characteristic wavenumber associated with the magnetic field as
$k_{\bm{B}}\equiv\frac{\left(\bm{\nabla}\times\bm{B}\right)\cdot\bm{B}}{B^{2}}~{}.$ (11)
In the weak-field regime ($B\ll B_{\mathrm{QED}}$), we find $k_{5}\propto k_{\bm{B}}B^{2}/B_{\mathrm{QED}}^{2}\ll k_{\bm{B}}$. In contrast, in the strong-field regime ($B>B_{\mathrm{QED}}$) $k_{5}\lesssim k_{\bm{B}}$. Thus, the scale $k_{5}$ is ultimately bounded by the characteristic wavenumber of the background magnetic field.
@@
$\frac{d\varepsilon_{em}}{dt}$ $=$ $-\sigma_{e}{\bm{E}^{2}}-\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}-\frac{c}{4\pi}\bm{\nabla}\cdot\left(\bm{E}\times\bm{B}\right),$
$\frac{d\varepsilon_{5}}{dt}$ $=$ $-\frac{1}{2}\mu_{5}n_{5}\Gamma_{f}+\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}~{}.$ (13)
Here, each term on the right-hand side plays a distinct role. The first term in both equations represents a sink: Joule dissipation for magnetic energy and spin-flip processes for chiral energy. The final term in the magnetic energy equation corresponds to the Poynting flux, which becomes a surface term upon volume integration. The $\bm{E}\cdot\bm{B}$ terms describe the conservative exchange of energy between the magnetic field and chirality.
@@
[…]
In the absence of the Hall drift term, $c\,\bm{E}=\eta\left(\bm{\nabla}\times\bm{B}-k_{5}\bm{B}\right)$, and the chiral correction can offset magnetic dissipation if ×$\nabla\times\bm{B}\approx k_{5}\bm{B}$.
To assess the impact of the CME on magnetic field evolution in NSs, we performed magneto-thermal simulations using an extended 3D finite-volume version of the MATINS  code [25, 26, 27]. This version solves the coupled induction and heat diffusion equations, consistently incorporating chiral and spin-flip terms arising from the chiral anomaly. To focus on the NS crust, we apply potential-field boundary conditions (current-free magnetosphere) at an outer numerical boundary located at $\rho=10^{10}\,\mathrm{g\,cm}^{-3}$, near the transition between the liquid envelope and the solid crust. At the crust–core interface, we impose perfect-conductor boundary conditions. The temperature-dependent electrical conductivity is computed at each point of the star using the codes from the IOFFE repository [^2] http://www.ioffe.ru/astro/conduct/ [28]. In MATINS, the NS background model can be constructed using various zero-temperature equation of states (EOSs) from the CompOSE online database [^3] https:/compose.obspm.fr/. For this study, we adopt the BSk24 EOS [29], assuming a canonical NS with mass $M=1.4\,M_{\odot}$, radius $R=12.4$ km, and crustal thickness of 0.86 km.
@@
As detailed in Section II, we initialize our simulations with a turbulent magnetic field concentrated at small scales, with an average strength of $\sim 10^{16}$ G [6, 33]. The field is initially divided between poloidal and toroidal components, with a dominant toroidal contribution, consistent with the small-scale energy spectra predicted by proto-NS dynamo simulations [6]. We limit the dipolar component to a few $10^{12}$ G, much weaker than the large-scale fields inferred observationally for magnetars. The total magnetic energy, of order a few $10^{49}$ erg, remains small compared to the star’s gravitational binding energy ($\sim 10^{53}$ erg) or the rotational energy of a millisecond pulsar ($\sim 10^{51-52}$ erg).
@@
To assess the impact of this residual asymmetry, Figure 1 shows the total magnetic energy spectra for Run D (left) and Run F (right) at $t=0$, $t=30$ yr, and $t=300$ yr. A significant energy transfer—about two orders of magnitude—occurs toward multipoles with initially weak fields ($\ell\leq 20$), as the emergence of a nonzero $k_{5}$ enables the CME to redistribute magnetic energy across spatial scales, boosting weaker regions at the expense of stronger ones. A particularly striking feature is the strong amplification of the dipolar component ($\ell=1$), marking the natural formation of the largest and hardest-to-form scale. In Run D, the initial slope (nearly linear in $\ell$) at $\ell\leq 20$ is preserved but amplified, whereas in Run F, the slope at large-scales evolves toward an approximately linear dependence on $\ell$, leading to comparable magnetic energy spectra in both runs. In both cases, the CME preserves the spectral peak at $\ell_{0}$, which – being at small scales – is more susceptible to Ohmic dissipation over tens of kiloyears. By contrast, an inverse cascade typically shifts the peak toward lower $\ell$, enhancing the longevity of the large-scale field [31].
While this process may resemble an inverse cascade, its underlying mechanism is fundamentally different. In the presence of a helical magnetic field (as in our configuration) and an active Hall term, an inverse cascade can occur, though its impact is typically much weaker [32]. In such cases, nonlinear interactions, as shown by Frisch [35], transfer energy and helicity to larger scales through mode couplings that satisfy $\bm{k}=\bm{p}+\bm{q}$ with $|\bm{k}|\leq\max(|\bm{p}|,|\bm{q}|)$. By contrast, the CME operates simultaneously on all multipoles, enabling the redistribution of magnetic energy from small to larger scales in our simulations
@@
Our research reveals that magnetic helicity, a classical attribute, triggers a subtle chiral asymmetry—a quantum anomaly—in macroscopic astrophysical bodies like NSs. This faint chirality significantly shapes magnetic field evolution over centuries, overcoming robust spin-flip suppression. The CME restructures the magnetic field spectrum by redistributing energy across scales, converging toward a nearly linear slope in multipole degree $\ell$, particularly for $\ell<10$, which governs large-scale astrophysical observables.
@@ Acknowledgements.
We gratefully acknowledge the generous hospitality and stimulating discussions with Sanjay Reddy and the participants of the programs “IReNA‐INT Joint Workshop on Thermal and Magnetic Evolution of Neutron Stars’’ at the University of Washington, Seattle, and “Breaking New Ground in Supernova Physics 2025: Crossroads of Turbulence, Chiral Dynamics, and Machine Learning” at Fukuoka University, Japan. CD acknowledges support from the Ministerio de Ciencia, Innovación y Universidades, co-funded by the Agencia Estatal de Investigación, the Unión Europea (FSE+), and the Universidad de Alicante. Her contract is part of the fellowship JDC2023-052227-I, funded by MCIU/AEI/10.13039/501100011033 and the FSE+. We also acknowledge support from the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana, through grant CIPROM/2022/13, and from the AEI grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033.
@@
In the following sections, we first decompose the induction equation (excluding the Hall term) into its poloidal and toroidal magnetic field components to demonstrate how the chiral current induces coupling between them (Section I). We then describe the setup of the initial magnetic field used in this study (Section II), followed by an extended analysis of our simulation results (Section III).

### v2

@@ Abstract
The chiral magnetic effect (CME)—a macroscopic manifestation of the quantum chiral anomaly—induces currents along magnetic field lines, facilitating mutual conversion between chiral asymmetry and magnetic helicity. Although the finite electron mass suppresses chiral asymmetry through spin-flip processes, we demonstrate that the CME effectively shapes magnetar field evolution. Magnetic helicity acts as a persistent internal source of chiral asymmetry, which mediates the redistribution of magnetic energy across spatial scales, without requiring an external energy source. Our three-dimensional magneto-thermal simulations of the neutron star crust reveal a novel mechanism that significantly reconfigures the magnetic field inherited at birth, amplifying both toroidal and poloidal large-scale dipolar components (crucial for spin-down) to strengths of $10^{14}$ G within just a century, at the expense of small-scale structures. This astrophysical application of the CME, distinct and complementary to conventional hydrodynamic dynamo models, offers an innovative framework for understanding magnetar field dynamics and provides a transformative solution to the origin of their exceptionally strong, large-scale fields.
@@
Beyond macroscopic hydrodynamic processes such as dynamos and turbulence, quantum field theory establishes a fundamental microscopic connection between chirality and helicity through the chiral anomaly. This phenomenon facilitates a two-way transfer between fermionic chiral asymmetry and magnetic helicity. Recent research has increasingly focused on linking the generation of large-scale magnetic fields to chiral asymmetries produced during core-collapse supernovae and the proto-NS phase [8, 9]. However, significant challenges persist, most notably the efficiency of spin-flip scattering processes induced by the finite electron mass, raising doubts about the CME’s relevance in these environments (see Kamada et al. [10] for a review on the CME in different scenarios). The rapid, temperature-dependent spin-flip reactions reduce the chiral imbalance, constraining the effectiveness of the chiral instability mechanism [11, 12].
@@
NSs exhibit a complex internal structure with multiple fluid components, each possessing distinct hydrodynamical velocities. In the outer crust, a rigid ionic lattice severely limits ion mobility, allowing only electrons to flow freely and sustain the currents essential for magnetic field evolution. The inner crust introduces added complexity with superfluid neutrons, which partially decouple from the nuclear lattice and act as a neutral fluid component. Complexity peaks in the NS core, where coexisting superfluid neutrons and superconducting protons necessitate a sophisticated multi-fluid framework. Given this complexity, magnetic field evolution is typically studied using region-specific approximations. In this context, we focus on the crust, where only (slowly drifting) electrons are mobile within the solid ionic lattice, making the e-MHD approximation applicable [20]. The magnetic field evolution is then governed by the induction equation derived from Faraday’s law:
@@
Magnetic field evolution must be coupled to the evolution equation for the chiral number density $n_{5}\equiv n_{R}-n_{L}$ [17, 11, 21, 10], which includes both source and sink terms:
@@
$\frac{\partial(\bm{A}\cdot\bm{B})}{\partial t}$ $=$ $-2c\bm{E}\cdot\bm{B}-c\bm{\nabla}\cdot\left(\bm{E}\times\bm{A}\right).$ (5)
@@
Here, $Q_{5}=\int n_{5}\,dV$ is the total axial charge, $\chi_{m}=\int\bm{A}\cdot\bm{B}\,dV$ is the total magnetic helicity, and ${\Gamma_{5}}=\int n_{5}\Gamma_{f}dV$ is the average spin-flip rate. Total helicity is not strictly conserved due to the sink term $\Gamma_{5}$. We neglect both the helicity flux across the boundary $\left(\propto\bm{E}\times\bm{A}\right)$ and the weak interaction processes ($\Gamma_{w}^{\mathrm{eff}}\ll\Gamma_{f}$), as their contributions are marginal. Weak interactions could, however, transiently enhance the effect if the NS is out of chemical equilibrium. For simplicity, we set $\Gamma_{w}^{\mathrm{eff}}=0$ hereafter, since its typical values ($10^{-2}-10^{2}$ s-1) are many orders of magnitude smaller than the spin-flip rates (Eq. (7)).
@@
which is on the order of $10^{15}-10^{17}$ s-1 under typical neutron star conditions during the first few centuries. Since $\sigma_{e}$ depends on temperature, so does $\Gamma_{f}$, underscoring the need to consider the coupled evolution of temperature and magnetic field in simulations.
@@
$n_{5}(t)=\left(\frac{2\alpha}{\pi\hbar}\frac{\bm{E}\cdot\bm{B}}{\Gamma_{f}}\right)\left(1-e^{-\Gamma_{f}t}\right)+n_{5}^{0}e^{-\Gamma_{f}t},$ (8)
@@
$n_{5}(t)\approx\frac{2\alpha}{\pi\hbar}\frac{\bm{E}\cdot\bm{B}}{\Gamma_{f}}~.$ (9)
@@
Equation (10) highlights the two main contributions to the chiral asymmetry: the numerator originates from magnetic helicity and increases when the pointwise current helicity density, $\bm{B}\cdot(\bm{\nabla}\times\bm{B})$, grows – that is, when $\bm{\nabla}\times\bm{B}$ becomes increasingly aligned with $\bm{B}$ [26]. This alignment characterizes the so-called helical magnetic fields, where toroidal and poloidal components reinforce each other. In contrast, the spin-flip term in the denominator suppresses chiral asymmetry. If the magnetic field configuration possesses a non-zero average magnetic helicity, a small but appreciable chiral asymmetry can be sustained despite the action of spin-flip processes. In this regime, where a nonzero chiral asymmetry is sustained and $k_{5}$ evolves in quasi-equilibrium, one can insert Equation (10) into Equation (3), explicitly factoring out the chiral anomaly from the induction equation. Consequently, $k_{5}$ acts as a catalyst for energy transfer across scales, making the magnetic field evolution depend solely on the field itself. In this formulation, the chiral contribution manifests as a nonlinear correction associated with the current component parallel to $\bm{B}$, while the Hall term emerges as a quadratic nonlinearity in $\bm{B}$, governed by the current component perpendicular to the field.
Defining the characteristic wavenumber associated with the current parallel to the magnetic field as
$k_{\bm{B}}\equiv\frac{\left(\bm{\nabla}\times\bm{B}\right)\cdot\bm{B}}{B^{2}}~,$ (11)
we can write
$k_{5}=\frac{k_{\bm{B}}}{1+\left(\dfrac{2\mu_{e}^{2}}{m_{e}^{2}c^{4}}\right)\dfrac{B^{2}_{\mathrm{QED}}}{3\pi B^{2}}}.$ (12)
In the weak-field regime ($B\ll B_{\mathrm{QED}}$), we find $k_{5}\propto k_{\bm{B}}B^{2}/B_{\mathrm{QED}}^{2}\ll k_{\bm{B}}$. In the strong-field regime ($B>B_{\mathrm{QED}}$), $k_{5}$ is ultimately bounded by $k_{\bm{B}}$ ($k_{5}\lesssim k_{\bm{B}}$).
@@
$\frac{\partial\varepsilon_{em}}{\partial t}$ $=$ $-\sigma_{e}{\bm{E}^{2}}-\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}-\frac{c}{4\pi}\bm{\nabla}\cdot\left(\bm{E}\times\bm{B}\right),$
$\frac{\partial\varepsilon_{5}}{\partial t}$ $=$ $-\frac{1}{2}\mu_{5}n_{5}\Gamma_{f}+\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}~.$ (14)
Here, $\varepsilon_{em}=B^{2}/8\pi$ and $\varepsilon_{5}=\mu_{5}n_{5}/2$, with each term on the right-hand side plays a distinct role. The first term in both equations represents a sink: Joule dissipation for magnetic energy and spin-flip processes for chiral energy. The final term in the magnetic energy equation corresponds to the Poynting flux, which becomes a surface term upon volume integration. The $\bm{E}\cdot\bm{B}$ terms describe the conservative exchange of energy between the magnetic field and chirality.
@@
where $E_{em}$ and $E_{5}$ are the volume integrals of $\varepsilon_{em}$ and $\varepsilon_{5}$, respectively. $S_{\mathrm{tot}}=\dfrac{c}{4\pi}\oint dS\cdot\left(\bm{E}\times\bm{B}\right)$ is the total Poynting flux, ${\Gamma_{5}}_{\mathrm{tot}}=\frac{1}{2}\int\mu_{5}n_{5}\Gamma_{f}\,dV$ is the total spin-flip dissipation rate, and $Q_{\mathrm{tot}}$ is the total Joule dissipation, given by:
$Q_{\mathrm{tot}}=\int\sigma_{e}\bm{E}^{2}dV.$ (16)
In the absence of the Hall drift term, $c\,\bm{E}=\eta\left(\bm{\nabla}\times\bm{B}-k_{5}\bm{B}\right)$, and the chiral correction can offset magnetic dissipation if $\nabla\times\bm{B}\approx k_{5}\bm{B}$.
To assess the impact of the CME on magnetic field evolution in NSs, we performed magneto-thermal simulations using an extended 3D finite-volume version of the MATINS  code [27, 28, 29]. This version solves the coupled induction and heat diffusion equations, consistently incorporating chiral and spin-flip terms arising from the chiral anomaly. To focus on the NS crust, we apply potential-field boundary conditions (current-free magnetosphere) at an outer numerical boundary located at $\rho=10^{10}\,\mathrm{g\,cm}^{-3}$, near the transition between the liquid envelope and the solid crust. At the crust–core interface, we impose perfect-conductor boundary conditions. The temperature-dependent electrical conductivity is computed at each point of the star using the codes from the IOFFE repository [^2] http://www.ioffe.ru/astro/conduct/ [30]. In MATINS, the NS background model can be constructed using various zero-temperature equations of state (EOS) from the CompOSE online database [^3] https:/compose.obspm.fr/. For this study, we adopt the BSk24 EOS [31], assuming a canonical NS with mass $M=1.4\,M_{\odot}$, radius $R=12.4$ km, and crustal thickness of 0.86 km.
@@
As detailed in Section II of the Supplemental Material, we initialize our simulations with a turbulent, partially helical444 A single-mode helical field is force-free, with the current everywhere parallel to the magnetic field ($(\bm{\nabla}\times\bm{B})\parallel\bm{B}$), leading to a zero Lorentz force. However, combining modes with different wavenumbers generally does not produce a force-free condition due to mode interactions. magnetic field (see Appendix A of Reisenegger [35]) with $B\gg B_{\rm sat}$, concentrated at very small scales and with an average strength of $\sim 10^{16}$ G [6, 36]. The field is initially divided between poloidal and toroidal components, with a dominant toroidal contribution, consistent with the small-scale energy spectra predicted by proto-NS dynamo simulations [6]. We limit the dipolar component to a few $10^{12}$ G, much weaker than the large-scale fields inferred observationally for magnetars. The total magnetic energy, of order a few $10^{49}$ erg, remains small compared to the star’s gravitational binding energy ($\sim 10^{53}$ erg) or the rotational energy of a millisecond pulsar ($\sim 10^{51-52}$ erg).
@@
To assess the impact of this residual asymmetry, Figure 1 shows the total magnetic energy spectra for Run D (left) and Run F (right) at $t=0$, $t=30$ yr, and $t=300$ yr. A significant energy transfer—about two orders of magnitude—occurs toward multipoles with initially weak fields ($\ell\leq 20$), as the emergence of a nonzero $k_{5}$ enables the CME to redistribute magnetic energy across spatial scales, boosting weaker regions at the expense of stronger ones. A particularly striking feature is the strong amplification of the dipolar component ($\ell=1$), marking the natural formation of the largest and hardest-to-form scale. In Run D, the initial slope, nearly linear in $\ell$ for $\ell\leq 20$, is preserved but amplified. Conversely, in Run F, the slope at large scales gradually evolves toward a similar linear dependence on $\ell$. Ultimately, both runs yield comparable magnetic energy spectra, suggesting that each mode saturates at a distinct characteristic amplitude. In both cases, the CME preserves the spectral peak at $\ell_{0}$, which – being at small scales – is more susceptible to Ohmic dissipation over tens of kiloyears. By contrast, an inverse cascade typically shifts the peak toward lower $\ell$, enhancing the longevity of the large-scale field [33].
While this process may resemble an inverse cascade, its underlying mechanism is fundamentally different. In the presence of a helical magnetic field (as in our configuration) and an active Hall term, an inverse cascade can occur, though its impact is typically much weaker [34]. In such cases, nonlinear interactions—specifically those involving the component of the current density perpendicular to the magnetic field—transfer energy and helicity to larger scales through mode couplings that satisfy $\bm{k}=\bm{p}+\bm{q}$ with $|\bm{k}|\leq\max(|\bm{p}|,|\bm{q}|)$, as demonstrated by Frisch [38]. By contrast, the CME, which involves the component of the electric current parallel to the magnetic field, operates simultaneously and independently on all multipoles (see Section I of the Supplemental Material), enabling the redistribution of magnetic energy from small to larger scales in our simulations.
@@
Our research reveals that magnetic helicity, a classical attribute, triggers a subtle chiral asymmetry—a quantum anomaly—in macroscopic astrophysical bodies like NSs. This faint chirality significantly shapes magnetic field evolution over centuries, overcoming robust spin-flip suppression. The CME restructures the magnetic field spectrum by redistributing energy across scales, converging toward a nearly linear slope in multipole degree $\ell$, particularly for $\ell<10$, which governs large-scale astrophysical observables. This linearity arises because each mode saturates at its characteristic amplitude.
@@ Acknowledgements.
We thank anonymous referee B for a careful report that improved the manuscript. We gratefully acknowledge stimulating discussions with Axel Brandenburg, Sanjay Reddy, Andrei Beloborodov, Chris Thompson and Valentin Skoutnev and other participants at the “IReNA‐INT Joint Workshop on Thermal and Magnetic Evolution of Neutron Stars’’ (University of Washington), “Breaking New Ground in Supernova Physics 2025: Crossroads of Turbulence, Chiral Dynamics, and Machine Learning” (Fukuoka University), and “Extreme Physics of Neutron Star Interiors” (Princeton University). We thank the organizers for their hospitality and for inviting CD as a speaker. CD is supported by the Ministerio de Ciencia, Innovación y Universidades (JDC2023-052227-I), co-funded by AEI (MCIN/AEI/10.13039/501100011033), the FSE+, and the Universidad de Alicante. CD and JP acknowledge support from the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana, through grant CIPROM/2022/13, and from the AEI grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033.
@@
In the subsequent sections, we initially break down the induction equation (omitting the Hall term) into its poloidal and toroidal components to show how the chiral current links the different modes (Section I). Although Section I presents a simplified toy model, treating the chiral wavenumber $k_{5}$ and magnetic diffusivity $\eta$ as constants for a clear, quasi-analytical illustration of the proposed mechanism, the simulations themselves were performed using a full 3D code, without these simplifying assumptions. We then describe the setup of the initial magnetic field used in this study (Section II).